OpenStudy (anonymous):

simplify: -36/ [-4-(-8)]*3^2

OpenStudy (anonymous):

-36/(-4+8)(9)=-36/(4*9)=-36/36=-1

OpenStudy (anonymous):

can u help me with one more plz

OpenStudy (anonymous):

-81

OpenStudy (anonymous):

which one is it

OpenStudy (anonymous):

Refer to the attachment.

OpenStudy (anonymous):

thanks